Use the method of separation of variables to determine the general solution to the one dimensional heat equation 2u/∂x2=1/α(∂u/∂t),0≤ x l,t≥0 subject to the boundary conditions u(0,t)=u((l,t)=0 and the initial condition u(x,0)=f(x). Determine the particular solution when f(x)=x2

1
Expert's answer
2021-10-19T16:40:02-0400

"u(x,t)=X(x)T(t)"

"X(x)T'(t)=1/\alpha\cdot T(t)X''(x)"

"\alpha T'(t)/T(t)=X''(x)/X(x)=-\lambda"


"T'(t)=-\lambda T(t)/\alpha"

"X''(x)+\lambda X(x)=0"

"X(0)=0,X(l)=0"


"\lambda>0"

"X(x)=c_1cos(\sqrt{\lambda}x)+c_2sin(\sqrt{\lambda}x)"

"0=X(0)=c_1"

"0=X(l)=c_2sin(l\sqrt{\lambda})"

"sin(l\sqrt{\lambda})=0\implies l\sqrt{\lambda}=n\pi,\ n=1,2,3,..."

"\lambda_n=(\frac{n\pi}{l})^2,\ X_n(x)=sin(\frac{n\pi x}{l})"


"\lambda=0"

"X(x)=c_1+c_2x"

"0=X(0)=c_1,\ 0=X(l)=c_2l\implies c_2=0"

So, in this case the only solution is the trivial solution and so λ=0 is not an eigenvalue for this boundary value problem.


"\lambda<0"

"X(x)=c_1cosh(\sqrt{-\lambda}x)+c_2sinh(\sqrt{-\lambda}x)"

"0=X(0)=c_1"

"0=X(l)=c_2sinh(l\sqrt{-\lambda})"

"\lambda_n=(\frac{n\pi}{l})^2,\ X_n(x)=sin(\frac{n\pi x}{l})"


for the time differential equation:

"T(t)=ce^{-k\lambda_nt}=ce^{-k(\frac{n\pi}{l})^2t}"


the general solution:

"u_n(x,t)=B_nsin(\frac{n\pi x}{l})e^{-k(\frac{n\pi}{l})^2t}"


where


"B_n=\frac{2}{l}\int^l_0 f(x)sin(\frac{n\pi x}{l})dx=\frac{2}{l}\int^l_0 x^2sin(\frac{n\pi x}{l})dx="


"=\frac{2}{n^3\pi^3}(2\pi lnxsin(n\pi x/l)+(2l^2-n^2\pi^2x^2)cos(n\pi x/l))|^l_0="


"=\frac{2}{n^3\pi^3}((2l^2-n^2\pi^2l^2)cos(n\pi )-2l^2)"


for odd n:

"u_n(x,t)=-\frac{2l^2}{n^3\pi^3}((4+n^2\pi^2)sin(\frac{n\pi x}{l})e^{-k(\frac{n\pi}{l})^2t}"


for even n:

"u_n(x,t)=-\frac{2l^2}{n\pi}"